OpenStudy (anonymous):

Partial fraction decomposition (s-4)/(s^2+4)(s+1)

OpenStudy (amistre64):

(s-4) A Bx+C ----------- = ---- + ------ (s^2+4)(s+1) s+1 s^2+4

myininaya (myininaya):

s=x

OpenStudy (anonymous):

(s-4)/(s^2 +4)(s+1) = As+B/(s^2+4) + C/(s+1) Multiply both sides by the denominator s-4 = (As+B)(s+1) + C(s^2+4) now plug numbers in. Plug in s=-1 -5 = 0 + 5C C = -1 s=0 s-4 = (As+B)(s+1) - (s^2+4) -4 = B - 4 B = 0 s-4 = As(s+1) - (s^2 + 4) plug in 1 -3 = 2A - 5 2 A = 2 A = 1 so A = 1 B = 0 C = -1 Back to the original... As+B/(s^2+4) + C/(s+1) s/(s^2+4) - 1/(s+1)

OpenStudy (amistre64):

(s-4) = A(s^2+4) + (Bx+C)(s+1) when s = -1 we get -5 = A(s^2+4) + (Bx+C)0 -5 = A(5) A = -1

myininaya (myininaya):

lol amistre you keep writing x

OpenStudy (amistre64):

(s-4) = -(s^2+4) + (Bx+C)(s+1) (s-4) = -s^2-4 + Bx(s+1)+ C(s+1); this can get tricker :)

OpenStudy (amistre64):

lol ... bummer

OpenStudy (amistre64):

s-4= -s^2 -4 +B(s^2) +B(s) + C(s) + C

OpenStudy (amistre64):

now we relate variables to coefficients i think

OpenStudy (amistre64):

s^1 +s = (B)s^2 +(B+C)s +C

OpenStudy (amistre64):

s^2 +s = (B)s^2 +(B+C)s +C B = 1; C = 0

OpenStudy (amistre64):

(s-4) -1 s ----------- = ----- + ------ ; maybe :) (s^2+4)(s+1) s+1 s^2+4

myininaya (myininaya):

A+B=0 => A=-B B+C=1 => C=1-B 4A+C=-4 => 4(-B)+(1-B)=-4 => -5B+1=-4 B=1 C=1-1=0 A=-1

myininaya (myininaya):

thats right! gj

OpenStudy (anonymous):

\[= A/(s+1) + (Bs + C)/(s ^{2}+ 4).\] We can clear this equation of fractions to get s-4=\[ s - 4 = A(s ^{2}+4) + (Bs + C)(s+1).\]

OpenStudy (amistre64):

\frac{top}{bottom} gives you a nice fraction \[\frac{top}{bottom}\]

OpenStudy (anonymous):

(A+ B)s^2 + (4A + B + C) s + (B + 4C) = s - 4 A + B = 0 4A + B + C = 1 B + 4C = -4

OpenStudy (anonymous):

I may have miscalculated somewhere, so TIWAGOS. The method I described here will always work for partial fractions decompositions.